complete the problem using the appropriate mathmatical teq. even though the answer you get will not be a whole number, it should be used to make your recommendation as to the make- up of the committee.
The president of a company would like to create a committee of employees to discuss issues relating to employee morale and productivity. The committee will then make policy recommendations to management. The president would like the make up of the committee to mirror the make up of the company in terms of the numbers men and women. If the ratio of men to women in the company is 84 to 147 how many of each gender should be selected to form a committee of 7?

Answer by stanbon(75887)

If the ratio of men to women in the company is 84 to 147 how many of each gender should be selected to form a committee of 7?
----------
84/147 = x/7
x=7(84/147)
x=4
Four women should be selected.
Cheers,
Stan H.